# Large Language Models in Argument Mining: A Survey

### v5

@@ 7.5 Theory-Grounded Modelling
Most LLM pipelines treat argument quality as a monolithic scalar. Argumentation theory, however, differentiates logical sufficiency, dialectical relevance and rhetorical effectiveness. Embedding these distinctions into prompts, loss functions or evaluation metrics is still nascent. Early attempts to weave Walton’s critical questions into dialogue evaluation improved error localisation but slowed inference (Palmieri, 2024). Future work could explore neuro-symbolic hybrids that use graph-based logic checkers as fast filters before LLM re-drafting.
7.6 Human-Centred and Interactive AM
Educational feedback tools, civic debate platforms and policy consultation interfaces require models that operate with users, not merely for them. That includes tracking my-side bias (Baccini and Hartmann, 2022), adapting critique to a learner’s knowledge state (Diana et al., 2020), and providing actionable, value-aligned suggestions. Reinforcement learning from argumentative feedback, real-time uncertainty estimates and personalisation via retrieval of prior user arguments are promising building blocks.

### v6

@@ 7.5 Theory-Grounded Modelling
Current LLM pipelines often operationalise argument quality as a single scalar, thereby flattening distinctions central to argumentation theory—logical sufficiency, dialectical relevance, and rhetorical effectiveness. Incorporating these dimensions into prompts, training objectives, and evaluation metrics remains underexplored. Early attempts to inject Walton’s critical questions into evaluation protocols improve error localisation but slow inference (DBLP:conf/cmna/Palmieri24). Future research may benefit from neuro-symbolic architectures in which structured logic checkers act as fast filters or constraint modules that prune implausible argumentative steps before generation. Progress in this direction would help align computational models more closely with the principles of formal and informal argumentation.
7.6 Human-Centred and Interactive Argument Mining
Argument mining increasingly serves interactive, user-facing applications: educational feedback tools, civic debate platforms, and policy consultation interfaces. These settings require systems that work with users—adapting to their goals, preferences, and knowledge states—rather than merely predicting labels. Relevant capabilities include tracking my-side bias (DBLP:conf/cogsci/Baccini022), tailoring critiques to a learner’s expertise (DBLP:conf/chi/DianaSK20), and providing actionable, value-aligned suggestions. Reinforcement learning from argumentative feedback, retrieval of users’ past arguments, and real-time uncertainty estimation could form the basis for next-generation interactive AM systems. Realising this vision will require collaborations across machine learning, argumentation theory, cognitive science, and HCI.